Short Report – QA/QC Meeting II Mar 2022
Meeting Date: 31 Mar 2022
- Checking recent fluxes from our stations
- Some site scripts on the RDS need an update, please contact LH. Still open: CH-CHA, CH-OE2
- Updated dataset CH-DAV FP2021.2: There was an issue with VPD on some days in July 2020, is now fixed. This changes 2020 gap-filled fluxes slightly, all other years are unaffected.
- Update on the database (db):
- The database is running without any issues so far, very stable.
- More variables of currently incoming data are added regularly. Soon all current variables are ingested.
- There will be meetings for showing how the database works, first meeting will be in May.
- If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
- SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)
- The purpose of QA/QC meetings is to check on current, incoming data. SRPs choose specific issues we should look at together and discuss in the group. Fluxes are checked if the respective SRP wishes to do so.
- There is a list of known variable abbreviations that you can use in case you wonder what an abbreviation means: Variable Abbreviations
- Check of EC raw data files:
- Recommended check for SRPs and T-SRPS: take a look at EC raw data files and check if they look OK
- Current EC raw data files are automatically converted to ASCII each day (done by the Python script bico)
- Files and their plots can be found here, e.g. for CH-LAS:
- Weekly flux calculations on the RDS:
- Please calculate fluxes and check them once per week, or more often if you wish to do so.
- If you cannot calculate the fluxes, try to find a substitute, e.g. LH.
- Please move your Level-0 results from the RDS to the respective Level-0 folder.
- Automatic upload of Level-0 fluxes to the database will be implemented soon.
- List of QA/QC Meeting dates: QA/QC Meetings 2022
LH, AS, IF, RM, LK, GA, RG, LS, TB, SB (10)
Peak Sahara dust in the air: 15 Mar 2022
Short data gap due to software issues.
Very clear and consistent CO2 uptake since first day of the year. Fertilization on 14.3., rain on 15.3. (only rain event in March). N2O and CH4 fluxes nearly complete, only one 24h-gap so far this year.
EC system runs well, but there was an issue with one of the sonic transducers (it popped out of the housing, found 9 Feb 2022, transducer was pushed back in), no issue visible in data. Sonic was replaced on 8 Mar 2022 (HS-100). Forest still covered in snow.
Net emission of CO2, forest is still covered in snow.
CO2 uptake started recently.
Unknown data format between mid-December and start of February, we will try find out the format. Data gap in February due to calibration. Recent fluxes look OK.
Subcanopy now shows CO2 uptake.
Already very clear CO2 uptake during the day since start of year.